OpenStudy (anonymous):

The sum of three numbers is 6, the first number is the sum of the second and third number. The second number is two more than the first.

OpenStudy (johnweldon1993):

x + y + z = 6 x = y + z y = x + 2 3 equations...3 unknowns....sounds Doable

OpenStudy (anonymous):

thank you very much